# Matchings avoiding ordered patterns

### v1

@@ 1 Introduction
In recent years, there has been accelerating interest on Turán-type questions for ordered graphs, see e.g. [17]. For a family of unordered graphs $\mathcal{H}$, the Turán number $ex({\mathcal{H}},n)$ denotes the maximum number of edges of an $n$-vertex graph, which does not contain a copy of any $H\in{\mathcal{H}}$. The celebrated Erdős–Stone–Simonovits theorem determines $ex({\mathcal{H}},n)$ asymptotically for $\mathcal{H}$ fixed, provided $\mathcal{H}$ contains no bipartite graph. Analogous notions exist for ordered graphs. We say an ordered graph $G$ contains a copy of an ordered graph $H$ if there exists an injective graph isomorphism from $H$ to $G$, which preserves the order of the vertices.
@@
Figure 1: From right to left: separated, nested and crossing pairs.
@@
Non-Separated $\frac{3(k-1)n}{2}-\frac{(2k-1-\{n+1\}_{k})(k-1+\{n+1\}_{k})}{2}$ New, see Theorem 2.1
@@ 1.1 Previous work on ordered matchings
For every $n,k\in\mathbb{N}$ with $n\geq 2k$ we have
@@
For every $n,k\in\mathbb{N}$ with $n\geq 2k$ we have
@@ 1.2 Organisation of the paper and notation
The remainder of the paper is organised as follow. In Section 2.1, we state and prove our new Turán-type results. These include Turán-type results for non-separated and non-nested matchings. We also prove exact Turán-type results for alternating path and $ex_{<}(n,\text{Cross-}\mathcal{M}_{k},\text{Sep-}\mathcal{M}_{k})$. In Section 3, we show how our Turán-type results for alternating paths and non-nested matchings yield consequences for Ramsey-type questions. Section 4 includes some final remarks and future research directions.
Notation. Suppose that $G$ is a graph with vertex set $[n]$. For any $I\subseteq[n]$, $G[I]$ denotes the subgraph of $G$ induced by the vertices/points of $I$. Let $E(X,Y)$ denote the edges between vertices $X$ and $Y$, and $e(X,Y)$ the number of these edges.
@@ 2.1 The Turán number of non-separated matchings
For every $n,k\in\mathbb{N}$ with $n\geq 2k-1$, we have
$ex_{<}(n,\neg\text{Sep-}\mathcal{M}_{k})=\left\lceil\frac{n-2k+1}{k}\right\rceil\left(\binom{2k-1}{2}-\binom{k-1}{2}\right)+\binom{2k-1-\{(n+1)\}_{k}}{2}$,
@@
We have $e(G_{t+1})=e(G_{t})+\binom{2k-1}{2}-\binom{k-1}{2}$. It is easy to check that, for $t_{0}=\lceil(n-2k+1)/k\rceil$, the graph $G_{t_{0}}$ has $n$ vertices and precisely the required number of edges. Suppose for a contradiction that $G_{t}$ contains a non-separated $k$-matching $M$, and take $t$ to be minimal with respect to this property. In particular, we have $t\geq 1$. If $M$ contains an edge lying in $V(G_{t+1})\setminus V(G_{t})$, then such an edge is separated to any edge in $V(G_{t})$. In that case, $M$ must lie entirely in $V(G_{t+1})\setminus V(G_{t})$, which is not possible as this is a set of only $k$ vertices. By minimality, $M\cap E(G_{t+1})$ is non-empty and, by the previous observation, all such edges must be incident to the last k1k-1 vertices of $G_{t}$. As $M$ contains no separated edges, this implies all edges of $M$ intersect the last k1k-1 vertices of $G_{t}$, a contradiction.
@@
Let $G$ be an $n$-vertex graph as in Claim 2.2. Next, we show $G$ contains a minimal non-edge of appropriate length.
@@
There exists some $x\in[n]$ such that $x(x+k)$ is a missing edge and the graph
$G[\{x+1,\dots,x+k-1\}]$ is complete.
Proof of the claim. Let $G$ be maximal with respect to adding edges while keeping the property that there is no non-separated $k$-matching. Let $uv$ be a missing edge, say $u<v$, which minimises vuv-u. In particular, all edges, except $uv$ must be present in $G[u,u+1,\dots,v]$. By maximality of $G$, it follows that $G+uv$ contains a non-separated $k$-matching $\mathcal{M}$, that contains the edge $uv$. Suppose that an edge $ab$ of the matching $\mathcal{M}$ lies below $uv$, say $u<a<b<v$. Now $\mathcal{M}\cup\{ub,av\}\setminus\{ab\}$ is a non-separated $k$-matching in $G$, contradiction. Suppose that there is a vertex $w$ with $u<w<v$ which is not saturated by the matching $\mathcal{M}$. Let $\mathcal{M}_{1}:=\mathcal{M}\cup\{uw\}\setminus\{uv\}$ and $\mathcal{M}_{2}:=\mathcal{M}\cup\{wv\}\setminus\{uv\}$. Both $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ are $k$-matchings in $G$. Therefore, by assumption, both of them contain a separated pair of edges. In particular, there exists an edge $e\in\mathcal{M}\setminus\{uv\}$ such that $e$ and $uw$ are separated. If $e$ lies to the left of $uw$, then $e$ is also separated to $uv$, contradicting the assumption that $\mathcal{M}$ is non-separated. Hence, $e$ lies to the right of $w$. Similarly, there exists an edge $e^{\prime}\in\mathcal{M}\setminus\{uv\}$, which lies to the left of $w$. Again, this is a contradiction, since $\mathcal{M}$ was non-separated.
@@
We prove Theorem 2.1 by induction on $n$. Let $G_{1}:=G[1,2,\dots,x{+}k-1]$ and $G_{2}:=G[x{+}1,\dots,n]$, where $x$ satisfies the conditions of Claim 2.3. By the inductive hypothesis, we have the following upper bounds on $e(G_{1})$ and $e(G_{2})$ (where we also use the equality $\lceil N/k\rceil=(N+\{N\}_{k})/k$ for any positive integer $N$).
$e(G_{1})\leq\frac{|G_{1}|-2k+1+\{|G_{1}|+1\}_{k}}{k}\left(\binom{2k-1}{2}-\binom{k-1}{2}\right)+\binom{2k-1-\{(|G_{1}|+1)\}_{k}}{2}$
$e(G_{2})\leq\frac{|G_{2}|-2k+1+\{|G_{2}|+1\}_{k}}{k}\left(\binom{2k-1}{2}-\binom{k-1}{2}\right)+\binom{2k-1-\{(|G_{2}|+1)\}_{k}}{2}$
@@
since $G_{1}\cap G_{2}$ is complete, and e([x],[x+k,n])=0e([x],[x{+}k,n])=0 by Claim 2.3 (since $x(x+k)$ is missing, there are no edges above it). Combining these equalities with the bounds on $e(G_{1})$ and $e(G_{2})$, we obtain
@@
Since $|G_{1}|+|G_{2}|=n+k-1$, we either have $\{|G_{1}|+1\}_{k}+\{|G_{2}|+1\}_{k}=\{n+1\}_{k}$ or $\{|G_{1}|+1\}_{k}+\{|G_{2}|+1\}_{k}=\{n+1\}_{k}+k$. In the former case, we have $f(\{|G_{1}|+1\}_{k},\{|G_{2}|+1\}_{k})\leq f(0,\{n+1\}_{k})$ by Claim 2.4. Combining this with inequality (1), we obtain
@@ 2.2 The Turán number of (strongly) non-nested matchings
There are many graphs achieving the lower bound in Theorem 2.5. Indeed, given an ordered graph $G$ containing no non-nested $k$-matching and with e(G)=(k1)|G|e(G)=(k-1)|G|, consider the following two operations.
@@
In the next three subsections, we provide more details on the behaviour of $ex_{<}(n,\neg\text{Nest-}\mathcal{M}_{k})$. In the first subsection, we prove the upper bound of Theorem 2.5. In the second subsection, we determine the maximum number of edges in a graph with no separated or crossing matching. In the third subsection, we propose a conjecture on the behaviour of $ex_{<}(n,\neg\text{Nest-}\mathcal{M}_{k})$ and discuss some implications on the analogous Ramsey problem.
@@
The lower bound follows from the Construction given earlier in the section. For the upper bound, it suffices to show that if $G$ is a graph with $V(G)=[n]$ and $G$ does not contain a non-nested $k$-matching, then $e(G)\leq(k-1)n+\binom{k-1}{2}$.
@@ 2.3 The Turán number of crossing versus separated matchings
For every $n,k\in\mathbb{N}$ with $n\geq 2k$ we have $ex_{<}(n,\text{Cross-}\mathcal{M}_{k},\text{Sep-}\mathcal{M}_{k})\leq(k-1)n-\binom{2k-1}{2}+\binom{k}{2}+(k-1)(\binom{k+1}{2}-1)$. Furthermore, equality holds provided $n\geq 2k^{2}$.
@@
It suffices to show that if $G$ is a graph with $V(G)=[n]$ and $G$ does not contain any crossing or separated $k$-matching, then $e(G)\leq(k-1)n-\binom{2k-1}{2}+\binom{k}{2}+(k-1)(\binom{k+1}{2}-1)$.
@@
Now suppose that $n\geq 2k^{2}$. We define a graph $G$ on vertex set $[n]$. Let $v_{i}=ik$ for $i=1,2,\ldots,k-1$ and let
@@
Now $e(G)=(k-1)n-\binom{2k-1}{2}+\binom{k}{2}+(k-1)(\binom{k+1}{2}-1)$.
We conjecture that the lower bound in Theorem 2.5 is optimal for strongly non-nested matching, and thus for non-nested matching.
Conjecture 2.9.
For every $n,k\in\mathbb{N}$ with $n\geq 2k$, we have
@@ 2.4 The Turán number of alternating paths
$ex_{<}(n,\text{Nest-}\mathcal{M}_{k})=\binom{n}{2}-\binom{n-2k+2}{2}$.
@@
Let $H$ be a graph with vertex set $[n]$, where $xy$ is an edge if and only if $|y-x|\leq 2k-2$. Given a nested $k$-matching on $[n]$, note that the edge $uv$ containing all other edges of the matching must satisfy $|u-v|\geq 2k-1$. Therefore, $H$ does not contain a nested $k$-matching. Since an alternating $2k$-path contains a nested $k$-matching, we conclude that $H$ does not contain an alternating $2k$-path either. It is easy to check that $e(H)=\binom{n}{2}-\binom{n-2k+2}{2}$, hence we have $ex(n,\text{Alt-}P_{2k})\geq\binom{n}{2}-\binom{n-2k+2}{2}$. It remains to show that $ex_{<}(n,\text{Alt-}P_{2k})\leq\binom{n}{2}-\binom{n-2k+2}{2}$.
Let $G$ be an arbitrary graph with vertex set $[n]$, which does not contain a copy of $\text{Alt-}P_{2k}$ and with $e(G)=ex_{<}(n,\text{Alt-}P_{2k})$. We define graphs $G_{0},G_{1},G_{2}\dots$ recursively as follows. We start with $G_{0}:=G$. Given $G_{i}$, for every $v\in[n]$ we pick $v^{-}$ and $v^{+}$ so that $v^{-}<v<v^{+}$, both $v^{-}v$ and $vv^{+}$ are edges in $G_{i}$ and, subject to this,
  • if $i$ is even, pick $v^{+}$ as large as possible and $v^{-}$ as small as possible,
  • if $i$ is odd, pick $v^{+}$ as small as possible and $v^{-}$ as large as possible.
Note that the vertices $v^{-}$ and $v^{+}$ may not exist e.g., if $v$ is isolated. We let $G_{i+1}$ be the graph obtained by deleting the edges $v^{-}v$ and $vv^{+}$ from $G_{i}$ for every $v\in[n]$.
Suppose that there exists an edge $x_{0}y_{0}$ in $G_{k-1}$ and that $k$ is even. By construction, it follows that there exists $x_{1}\in[y_{0}+1,n]$ such that $x_{0}x_{1}$ is an edge in $G_{k-2}$. Again, this implies there exists $x_{2}\in[1,x_{0}-1]$ such that $x_{2}x_{1}$ is an edge in $G_{k-3}$. Repeatedly applying this argument yields vertices $x_{1},\dots,x_{k-1}$, where $y_{0}<x_{1}<x_{3}<x_{5}<\dots<x_{k-1}$ and $x_{0}>x_{2}>x_{4}>\dots>x_{k-2}$ such that $x_{i}x_{i+1}$ is an edge in $G$ for every $i\in[k-1]\cup\{0\}$.
An analogous argument yields vertices $y_{1},\dots,y_{k-1}$, where $y_{i}y_{i+1}$ is an edge in $G$ for every $i\in[k-1]\cup\{0\}$, and $x_{0}<y_{1}<y_{3}<\dots<y_{4}<y_{2}<y_{0}$. Note that $y_{k-1}\dots y_{1}y_{0}x_{0}x_{1}\dots x_{k-1}$ is an alternating $2k$-path, a contradiction. The case were $k$ is odd is analogous.
We conclude that the graph $G_{k-1}$ must not have any edge. It follows that, for each vertex $v$ in $G$, there are at most 2k22k-2 integers $v^{\prime}>v$ such that $vv^{\prime}$ is an edge, as we delete at most $2$ such edges at each step. Similarly, there are at most 2k22k-2 integers $v^{\prime}<v$ such that $v^{\prime}v$ is an edge. This implies $ex_{<}(n,\text{Alt-}P_{2k})\leq\binom{n}{2}-\binom{n-2k+2}{2}$, as required. ∎

### v2

@@ 1 Introduction
In recent years, there has been accelerating interest on Turán-type questions for ordered graphs, see e.g. [17]. For a family of unordered graphs $\mathcal{H}$, the Turán number $ex(n,{\mathcal{H}})$ denotes the maximum number of edges of an $n$-vertex graph, which does not contain a copy of any $H\in{\mathcal{H}}$. The celebrated Erdős–Stone–Simonovits theorem determines $ex({n,\mathcal{H}})$ asymptotically for $\mathcal{H}$ fixed, provided $\mathcal{H}$ contains no bipartite graph. Analogous notions exist for ordered graphs. We say an ordered graph $G$ contains a copy of an ordered graph $H$ if there exists an injective graph isomorphism from $H$ to $G$, which preserves the order of the vertices.
@@
Figure 1: From left to right: separated, nested and crossing pairs.
@@
Non-Separated $\frac{3}{2}(k-1)n-\frac{1}{2}(2k-1-\{n+1\}_{k})(k-1+\{n+1\}_{k})$ New, see Theorem 2.1
@@ 1.1 Previous work on ordered matchings
For every $n,k\in\mathbb{N}$ with $n\geq 2k$, we have
@@
For every $n,k\in\mathbb{N}$ with $n\geq 2k$, we have
@@ 1.2 Organisation of the paper and notation
The remainder of the paper is organised as follow. In Section 2, we state and prove our new Turán-type results. These include Turán-type results for non-separated and non-nested matchings. We also prove exact Turán-type results for alternating path and $ex_{<}(n,\text{Cross-}\mathcal{M}_{k},\text{Sep-}\mathcal{M}_{k})$. In Section 3, we show how our Turán-type results for alternating paths and non-nested matchings yield consequences for Ramsey-type questions. Section 4 includes some final remarks and future research directions.
Notation. Suppose that $G$ is a graph with vertex set $[n]$. We write $e(G)$ for the number of edges in $G$. For any $I\subseteq[n]$, $G[I]$ denotes the subgraph of $G$ induced by the vertices/points of $I$. Let $E(X,Y)$ denote the edges between vertices $X$ and $Y$, and $e(X,Y)$ the number of these edges.
@@ 2.1 The Turán number of non-separated matchings
For every $n,k\in\mathbb{N}$ with $n\geq k$, we have
$ex_{<}(n,\neg\text{Sep-}\mathcal{M}_{k})=\left\lceil\frac{n-2k+1}{k}\right\rceil\left(\binom{2k-1}{2}-\binom{k-1}{2}\right)+\binom{2k-1-\{n+1\}_{k}}{2}$,
@@
We have $e(G_{t+1})=e(G_{t})+\binom{2k-1}{2}-\binom{k-1}{2}$. It is easy to check that, for $t_{0}=\lceil(n-2k+1)/k\rceil$, the graph $G_{t_{0}}$ has $n$ vertices and precisely the required number of edges. Suppose for a contradiction that $G_{t}$ contains a non-separated $k$-matching $M$, and take $t$ to be minimal with respect to this property. In particular, we have $t\geq 1$. If $M$ contains an edge lying in $V(G_{t+1})\setminus V(G_{t})$, then such an edge is separated to any edge in $V(G_{t})$. In that case, $M$ must lie entirely in $V(G_{t+1})\setminus V(G_{t})$, which is not possible as this is a set of only $k$ vertices. By minimality, $M\cap V(G_{t+1})$ is non-empty and, by the previous observation, all such edges must be incident to the last k1k-1 vertices of $G_{t}$. As $M$ contains no separated edges, this implies all edges of $M$ intersect the last k1k-1 vertices of $G_{t}$, a contradiction.
@@
In the next claim, we prove that a graph attaining $ex_{<}(n,\neg\text{Sep-}\mathcal{M}_{k})$ must contain a minimal missing edge of length approximately $k$. Note that this is also a property shared with the extremal construction (see the missing edge between consecutive blue points in Figure 2e).
@@
Let $G$ be a graph with $V(G)=[n]$ and $e(G)=ex_{<}(n,\neg\text{Sep-}\mathcal{M}_{k})$ which does not contain a non-separated $k$-matching. If $n\geq 2k$, there exists some $x\in[n]$ such that $x(x+k)$ is a missing edge and the graph $G[\{x+1,\dots,x+k-1\}]$ is complete.
Proof of the claim. By the assumptions, we known $G$ must have at least one missing edge and adding any missing edge must create a non-separated $k$-matching. Let $uv$ be a missing edge, say $u<v$, which minimises vuv-u. In particular, all edges, except $uv$ must be present in $G[u,u+1,\dots,v]$. By maximality of $G$, it follows that $G+uv$ contains a non-separated $k$-matching $\mathcal{M}$, that contains the edge $uv$. Suppose that an edge $ab$ of the matching $\mathcal{M}$ lies below $uv$, say $u<a<b<v$. Now $\mathcal{M}\cup\{ub,av\}\setminus\{ab\}$ is a non-separated $k$-matching in $G$, contradiction. Suppose that there is a vertex $w$ with $u<w<v$ which is not saturated by the matching $\mathcal{M}$. Let $\mathcal{M}_{1}:=\mathcal{M}\cup\{uw\}\setminus\{uv\}$ and $\mathcal{M}_{2}:=\mathcal{M}\cup\{wv\}\setminus\{uv\}$. Both $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ are $k$-matchings in $G$. Therefore, by assumption, both of them contain a separated pair of edges. In particular, there exists an edge $e\in\mathcal{M}\setminus\{uv\}$ such that $e$ and $uw$ are separated. If $e$ lies to the left of $uw$, then $e$ is also separated to $uv$, contradicting the assumption that $\mathcal{M}$ is non-separated. Hence, $e$ lies to the right of $w$. Similarly, there exists an edge $e^{\prime}\in\mathcal{M}\setminus\{uv\}$, which lies to the left of $w$. Again, this is a contradiction, since $\mathcal{M}$ was non-separated.
@@
We prove Theorem 2.1 by induction on $n$. The theorem trivially holds if $k\leq n\leq 2k-1$, since the extremal construction is simply the complete graph. This is the base step. Suppose instead that $n\geq 2k$. Let $G$ be a graph with $V(G)=[n]$ and $e(G)=ex_{<}(n,\neg\text{Sep-}\mathcal{M}_{k})$, which does not contain a non-separated $k$-matching and satisfied the properties of Claim 2.2. Let $x\in[n]$ satisfy the conditions of Claim 2.3 and define graphs $G_{1}:=G[1,2,\dots,x{+}k-1]$ and $G_{2}:=G[x{+}1,\dots,n]$. By the inductive hypothesis, since $G_{1}$ and $G_{2}$ have at least $k$ vertices each, we have the following upper bounds on $e(G_{1})$ and $e(G_{2})$ (where we also use the equality $\lceil N/k\rceil=(N+\{N\}_{k})/k$ for any positive integer $N$).
$e(G_{1})\leq\frac{|G_{1}|-2k+1+\{|G_{1}|+1\}_{k}}{k}\left(\binom{2k-1}{2}-\binom{k-1}{2}\right)+\binom{2k-1-\{|G_{1}|+1\}_{k}}{2}$
$e(G_{2})\leq\frac{|G_{2}|-2k+1+\{|G_{2}|+1\}_{k}}{k}\left(\binom{2k-1}{2}-\binom{k-1}{2}\right)+\binom{2k-1-\{|G_{2}|+1\}_{k}}{2}$
@@
since $G_{1}\cap G_{2}$ is complete (as $x(x+k)$ is minimal), and e([x],[x+k,n])=0e([x],[x{+}k,n])=0 by Claim 2.2 (as $x(x+k)$ is missing, there are no edges above it). Combining these equalities with the bounds on $e(G_{1})$ and $e(G_{2})$, we obtain
@@
Since $|G_{1}|+1+|G_{2}|+1=n+k+1$, we either have $\{|G_{1}|+1\}_{k}+\{|G_{2}|+1\}_{k}=\{n+1\}_{k}$ or $\{|G_{1}|+1\}_{k}+\{|G_{2}|+1\}_{k}=\{n+1\}_{k}+k$. In the former case, we have $f(\{|G_{1}|+1\}_{k},\{|G_{2}|+1\}_{k})\leq f(0,\{n+1\}_{k})$ by Claim 2.4. Combining this with inequality (1), we obtain
@@ 2.2 The Turán number of (strongly) non-nested matchings
In fact, we conjecture that the lower bound in Theorem 2.5 is optimal for strongly non-nested matching, and thus for non-nested matching.
Conjecture 2.6.
For every $n,k\in\mathbb{N}$ with $n\geq 2k$, we have
There are many graphs achieving the lower bound in Theorem 2.5 (and Conjecture 2.6). Indeed, given an ordered graph $G$ containing no non-nested $k$-matching and with e(G)=(k1)|G|e(G)=(k-1)|G|, consider the following two operations.
@@
The lower bound follows from the Construction given earlier in the section. For the upper bound, it suffices to show that if $G$ is a graph with $V(G)=[n]$ and $G$ does not contain a strongly non-nested $k$-matching, then $e(G)\leq(k-1)n+\binom{k-1}{2}$.
@@ 2.3 The Turán number of crossing versus separated matchings
For every $n,k\in\mathbb{N}$ with $n\geq 2k$ we have
$ex_{<}(n,\text{Cross-}\mathcal{M}_{k},\text{Sep-}\mathcal{M}_{k})\leq(k-1)n-\binom{2k-1}{2}+\binom{k}{2}+(k-1)\left(\binom{k+1}{2}-1\right).$
Furthermore, equality holds provided $n\geq 2k^{2}$.
@@
Firstly, we show that if $G$ is a graph with $V(G)=[n]$ and $G$ does not contain any crossing or separated $k$-matching, then $e(G)\leq(k-1)n-\binom{2k-1}{2}+\binom{k}{2}+(k-1)(\binom{k+1}{2}-1)$.
@@
To see that equality holds for $n\geq 2k^{2}$, consider the graph $G$ on vertex set $[n]$. Let $v_{i}=ik$ for $i=1,2,\ldots,k-1$ and let
@@
Now $e(G)=(k-1)n-\binom{2k-1}{2}+\binom{k}{2}+(k-1)(\binom{k+1}{2}-1)$. ∎
@@ 2.4 The Turán number of alternating paths
$ex_{<}(n,\text{Nest-}\mathcal{M}_{k})=2(k-1)n-(k-1)(2k-1).$
@@
Let $H$ be a graph with vertex set $[n]$, where $xy$ is an edge if and only if $|y-x|\leq 2k-2$. Given a nested $k$-matching on $[n]$, note that the edge $uv$ containing all other edges of the matching must satisfy $|u-v|\geq 2k-1$. Therefore, $H$ does not contain a nested $k$-matching. Since an alternating $2k$-path contains a nested $k$-matching, we conclude that $H$ does not contain an alternating $2k$-path either. It is easy to check that e(H)=2(k1)n(k1)(2k1)e(H)=2(k-1)n-(k-1)(2k-1), hence we have $ex_{<}(n,\text{Alt-}P_{2k})\geq 2(k-1)n-(k-1)(2k-1)$. It remains to show that $ex_{<}(n,\text{Alt-}P_{2k})\leq 2(k-1)n-(k-1)(2k-1)$.
Let $G$ be an arbitrary graph with vertex set $[n]$, which does not contain a copy of $\text{Alt-}P_{2k}$ and with $e(G)=ex_{<}(n,\text{Alt-}P_{2k})$. We define graphs $G_{0},G_{1},G_{2}\dots$ recursively as follows. We start with $G_{0}:=G$. Given $G_{i}$, for every $v\in[n]$ we pick $v^{+}\in[n]\setminus\{v\}$ such that $vv^{+}$is an edge in $G_{i}$ and
  • if $i$ is odd, $v^{+}>v$ and $v^{+}$ is as large as possible,
  • if $i$ is even, $v^{+}<v$ and $v^{+}$ is as small as possible.
Note that $v^{+}$ may not exist e.g., if $i$ is odd and $v$ has no neighbour on its right. We let $G_{i+1}$ be the graph obtained by deleting the edges $vv^{+}$ from $G_{i}$ for every $v\in[n]$.
Suppose that there exists an edge $x_{0}y_{0}$ in $G_{2k-2}$. By construction, it follows that there exists $x_{1}\in[y_{0}+1,n]$ such that $x_{0}x_{1}$ is an edge in $G_{2k-3}$. Again, this implies there exists $x_{2}\in[1,x_{0}-1]$ such that $x_{2}x_{1}$ is an edge in $G_{2k-4}$. Repeatedly applying this argument yields vertices $x_{1},\dots,x_{2k-2}$ where $\dots<x_{4}<x_{2}<x_{0}<y_{0}<x_{1}<x_{3}<x_{5}<\dots$ and $x_{i}x_{i+1}$ is an edge in $G$ for every $i\in[2k-2]$. But then $y_{0}x_{0}x_{1}\dots x_{2k-2}$ is an alternating $2k$-path in $G$, a contradiction.
We conclude that the graph $G_{2k-2}$ must not have any edge. Now, when we obtain $G_{i+1}$ from $G_{i}$, we delete at most one edge per vertex, namely the edge $vv^{+}$ for vertex $v$. We do not delete any edge when considering the largest non-isolated vertex (for $i$ odd) or the smallest non-isolated vertex (for $i$ even). Furthermore, it is easy to see that $G_{i}$ has at least $i$ isolated vertices (namely, $\{1,n,2,n-1,\dots\}$). We conclude that $e(G_{i+1})\geq e(G_{i})-(n-i-1)$ for every $i$. Hence, we have $0=e(G_{2k-2})\geq e(G_{0})-2n(k-1)-(k-1)(2k-1)$ and so $e(G)\leq 2(k-1)n-(k-1)(2k-1)$. This implies $ex_{<}(n,\text{Alt-}P_{2k})\leq 2(k-1)n-(k-1)(2k-1)$, as required. ∎